OpenStudy (anonymous):

The speed of light in a vacuum is 3.00 ´ 108 meters/second (m/s). There are 100 centimeters (cm) in 1 meter (m). How much time is needed for light to travel 1.00 cm in a vacuum

OpenStudy (ybarrap):

$$ \large{ 1~cm=.01m\\ \cfrac{1}{3(10)^8~m/s} \times .01~m\\ =(1/3)\times10^{-2}10^{-8}~s\\ =(1/3)\times10^{-10} } $$ Make sense?

OpenStudy (anonymous):

no, which one of these would be the answer? 3.33 mc015-1.jpg 10–11 s 3.33 mc015-2.jpg 10–9 s 3.00 mc015-3.jpg 1010 s 3.00 mc015-4.jpg 108 s

OpenStudy (ybarrap):

Do you know that speed = distance traveled / time taken? $$ speed=\cfrac{distance}{time}\\ \implies time=\cfrac{distance}{speed} $$ We know distance is .01 meters and speed is \(3(10)^8 m/s\). http://www.wolframalpha.com/input/?i=%281%2F3%2910%5E%28-10%29